A cell phone company orders 600 new phones from a manufacturer. If the probability of a phone being defective is 3.5%, predict how many of the phones are likely to be defective. Round to the nearest whole number.
Answer by ikleyn(52754)

.

3.5% of 600, which is  6*3.5 = 18 + 3 = 21.      ANSWER